# DSAI: Unbiased and Interpretable Latent Feature Extraction for Data-Centric AI

### v1

@@ 7 Conclusion
Appendix A Details of Datasets
To verify the applicability and effectiveness of our methodology, we utilized datasets from diverse industries, featuring a wide range of attributes and characteristics. Two datasets (slogan and research title) were used to validate our approach, while the other datasets were employed to demonstrate DSAI’s practical applications across various domains.
Slogan Dataset
We utilized the Slogan Generation Dataset Jin et al. (2023), a publicly available resource containing 5,014 validation examples, 1,322 test examples, and 251 human-annotated examples rated for "catchiness." Each entry includes a company name, a description of the company or product, the slogan, and its industry category. For example, the dataset features entries like Bikini Village Inc., categorized under "Apparel & Fashion," with the slogan "Your Brand Destination For Swimwear" and a description emphasizing its focus on trendy swimwear and beachwear. The dataset primarily features lesser-known companies, reducing bias from brand recognition and enabling analysis to focus on the intrinsic qualities of slogans across various industries.
Research Title Dataset
For analyzing research titles, we utilized a dataset of 10,874 research paper title and abstract pairs from previous work Wang et al. (2018). This dataset was preprocessed and divided into training (80%), validation (10%), and testing (10%) sets. On average, each title contains 9 words, and abstracts have 116 words.
MIND Dataset
The MIND dataset, created for news recommendation tasks, is a large-scale benchmark dataset sourced from Microsoft News Wu et al. (2020). It includes data from 1 million users and over 160,000 English news articles, each containing a title, abstract, and body text. We calculated click-through rates (CTR) based on the provided click and rejection logs for each article.
Spam Dataset
The SMS Spam Collection is a widely used dataset for spam detection research, containing 5,574 SMS messages in English, labeled as either "ham" (legitimate) or "spam" Kim (2016). This dataset was compiled from several public sources, including Grumbletext (425 manually extracted spam messages), the NUS SMS Corpus (3,375 legitimate messages from Singaporean students), Caroline Tag’s Thesis (450 ham messages), and the SMS Spam Corpus v.0.1 Big (1,002 ham messages and 322 spam messages).
Reddit Dataset
The Reddit dataset is a subset of a larger dataset from May 2019, containing 25,000 comments from each of the 40 most active subreddits Magnan (2019). Comments that were removed, deleted, or shorter than 4 tokens were excluded. The dataset includes features such as the subreddit where the comment was posted (categorical), the comment content (string), a binary controversiality metric, and a score indicating net upvotes minus downvotes.
Appendix B Expert-Defined Criteria
Previous research has devoted considerable effort into identifying the characteristics of "good content" in specific domains. Building upon this foundation, we conducted a preliminary investigation to establish expert-defined criteria for both slogans and research titles.
Slogan
To establish criteria for effective slogans, we synthesized insights from two foundational studies. The first study emphasized how slogans enhance brand identity, equity, and differentiation, providing guidelines such as ensuring long-term relevance, clear brand positioning, consistent brand association, creative depth for better recall, and minimal false recall Kohli et al. (2007). The second study outlined core characteristics of effective slogans, highlighting simplicity, pleasantness, compactness, persuasiveness, memorability, uniqueness, and honesty Padrakali and Chitra Chellam (2017). After thoroughly reviewing the two studies, we synthesized their insights into a unified set of expert-defined criteria as in Figure 11.
Title
For research titles, we similarly synthesized insights from two key studies Nair and Gibbert (2016); Tullu (2019). These studies offer comprehensive guidance on crafting impactful titles that align with academic expectations. They both emphasize that effective research titles should be clear, concise, and informative, accurately representing the study while capturing readers’ attention. Key recommendations include placing essential keywords early for indexing purposes, avoiding ambiguity and misleading content, and maintaining professional language. The studies stress the importance of balanced structure and linguistic clarity, with descriptive titles preferred over declarative or interrogative formats. We integrated these recommendations into a comprehensive set of expert-defined criteria as in Figure 11.
Appendix C LLM Annotation with Expert-Defined Criteria
@@
C.1 Method
We established a gold standard by manually annotating 10–20 samples per dataset and using expert-defined criteria. After optimizing prompts to maximize alignment with this manual annotation, we used the most aligned prompt to annotate 3,000 data points with GPT-4o, classifying each as "Conform" or "Not conform". The conformity score for each data point was calculated as the total number of conforming criteria. Based on these scores, we selected the top and bottom 600 samples as positive (high-quality text) and negative (low-quality text) samples respectively.
@@ C.1 Method
  Criterion Top Conform (%) Bottom Conform (%)
@@
Table 2: Ratio of Conform in Top and Bottom for Slogan expert-defined criteria.
C.2 Ensuring Distinction Between Positive and Negative Groups
Given that some expert criteria may lack sufficient discriminative power despite their theoretical importance, we conducted an analysis comparing the ratios of "Conform" for each criterion between top- and bottom-ranked samples. Note that a truly discriminative criterion should show high conformity in the top group and low conformity in the bottom group.
  Checklist Item Top Conform (%) Bottom Conform (%)
@@ C.2 Ensuring Distinction Between Positive and Negative Groups
Table 3: Ratio of Conform in Top and Bottom for Title expert-defined criteria.
@@
Table 2 demonstrate significant conformity ratio disparities between the top and bottom groups for most criteria. Several criteria demonstrate maximal distinction (e.g., "Concise but not too simple" (100% vs. 0%)), and some show moderate but still meaningful differentiation (e.g., "Include the brand name" (99% vs. 30%)). "No exaggeration" (99% vs. 91%) and "Direct/Straight-forward" (99% vs. 61%) have smaller gaps.
@@
Table 3 shows even stronger distinction between the top and bottom groups. Multiple criteria (e.g., "Simple format" "Direct", "Concise and precise") achieve 100% conformity in the top group compared to as low as 0–3% conformity in the bottom group, indicating their high predictive value for title quality. The large differences across all criteria confirm their strong discriminative power.
Appendix D Direct Latent Feature Extraction using LLMs
To evaluate whether the features directly generated by large language models (LLMs) are grounded in the input data or influenced by pre-existing knowledge, we conducted a series of experiments. These experiments were designed to investigate the effects of input data, label variations, and task context on feature generation (Section 3.2).
In these experiments, we ensured that all available data were utilized to maximize the representativeness of the input. For instance, we used all 600 positive samples for PosData and the complete set of 1200 samples (both positive and negative) for MixedData. This comprehensive use of data aimed to minimize bias and ensure that the features generated by the LLM were robustly grounded in the provided input.
To further address potential data-grounding issues in the direct feature generation process, we explicitly included instructions in the prompts emphasizing that the LLM must consider the input data when generating criteria. By reinforcing the importance of grounding, we sought to mitigate the risk of features being overly influenced by the model’s pre-existing knowledge, while the results revealed that the generated features were still significantly influenced by the LLM’s pre-existing knowledge (See Results in Section 3.2).
The specific prompts employed in each experimental setup are detailed in Figure 13, which provides a comprehensive outline of the instructions given to the LLM. The results of the experiments are presented in Figures 14 to 22, offering insights into the effectiveness of the approach and highlighting the model’s performance across various configurations.
Appendix E Implementation Details and Cost Analysis
In this section, we provide the prompts used in each stage of the DSAI pipeline along with corresponding outputs.
@@ Appendix E Implementation Details and Cost Analysis
In Stage #1 Perspective Generation, we generated 50 perspectives at per forwarding step across three steps, iteratively concatenating each step’s output with the few-shot example in the prompt for the next step. For Stage #2 Perspective-Value Matching, each forwarding step processed one sentence and three perspectives as input. Stage #3 Perspective-Oriented Value Clustering used all generated values per perspective in order to effectively cluster shared characteristics. In Stage #4 Verbalization, we transformed each perspective-label pair into a compact sentence (criterion). The prompts for all stages are in Figures 23 to 27.
@@
Appendix F Discriminative Power of Direct LLM- and DSAI-Generated Features
The accuracy with which generated features reflect the latent characteristics of the positive dataset depends on their alignment with the input data. To evaluate this alignment, we annotated the dataset using all generated features and assessed their discriminative power (DP) using two key metrics: $P(\textrm{positive}|\textrm{Conform})$ and $P(\textrm{negative}|\textrm{NotConform})$.
The metric $P(\textrm{positive}|\textrm{Conform})$ measures the proportion of positive data among instances labeled as "Conform" to a feature. This value indicates how effectively a feature captures the latent characteristics of the positive dataset. Scores below 0.5 suggest that the feature aligns more closely with negative data.
$P(\textrm{negative}|\textrm{NotConform})$ evaluates how successfully a feature identifies traits in the negative dataset. Scores above 0.5 indicate that the feature effectively captures the negative traits, while scores below 0.5 suggest misalignment, potentially indicating positive traits incorrectly marked as "Not Conform."
We used LLM labeling to filter out the "Not Applicable" cases (e.g., "technical terms in medical domains" for non-medical texts) before calculating DP scores.
This two-metric approach enables evaluation of features’ effectiveness in both capturing positive characteristics and identifying traits to avoid in negative data.
To simplify interpretation, in the following sections, we subtracted 0.5 from all DP scores. This adjustment highlights values below 0.5, making it easier to identify features that fail to align effectively with the intended dataset. Positive values after subtraction indicate alignment, while negative values signal misalignment.
F.1 Discriminative Power of LLM-Generated Features
Feature Slogan Title
PosData MixedData PosData MixedData
Feature 1 0.298535 0.061810 0.293475 0.214968
Feature 2 0.212195 0.193333 0.159765 0.075342
Feature 3 0.270186 0.295139 0.152727 0.049150
Feature 4 0.258410 0.031357 0.177354 0.169845
Feature 5 0.086310 0.069536 0.102517 0.160841
Feature 6 0.127063 0.120278 0.065525 0.170258
Feature 7 0.106327 0.227273 0.221347 0.191358
Feature 8 0.004625 0.128173 -0.116618 0.039169
Feature 9 0.333333 0.189394 0.163983 0.188144
Feature 10 -0.072835 0.004554 0.137737 0.170300
Feature 11 0.329457 0.023663 0.179561 0.141446
Feature 12 0.366667 0.151235 0.141446 0.219746
Feature 13 0.217842 0.005397 0.163901 -
Feature 14 0.288235 0.029469 0.105973 -
Feature 15 0.409091 0.008311 0.014444 -
Feature 16 0.343137 0.155172 0.105973 -
Feature 17 0.270492 0.269231 0.160693 -
Feature 18 0.203125 - 0.235832 -
Feature 19 - - 0.164412 -
Feature 20 - - 0.093699 -
Feature 21 - - -0.053386 -
Feature 22 - - 0.254810 -
Feature 23 - - -0.156301 -
Table 4: Features extracted from Slogan and Title datasets by PosData and MixedData. Values are subtracted by 0.5 to see values under 0.5 easily.
As shown in Table 4, several features from PosData and MixedData lacked the discriminative power and validity required for reliable analysis. In slogans, 3 out of 18 PosData features had DP scores below 0.1 (with 1 negative), and 8 out of 17 MixedData features were below 0.1. Similarly, in titles, 4 out of 23 PosData features were under 0.1 (with 3 negative), and 2 out of 13 MixedData features were under 0.1.
These results highlight that directly generating features from LLMs often produces features that lack sufficiently discriminative strength and reliability for robust data analysis, suggesting the need for improved methodologies.
  Category Bucket Feature 1 Feature 2 Feature 3 Feature 4 Feature 5 Feature 6 Feature 7 Feature 8 Feature 9 Average
  Slogan Bucket 1 0.3000 0.2969 0.1988 0.1791 0.1647 0.1446 0.1429 0.1114 0.0071 0.1644
Bucket 2 0.2963 0.2445 0.2442 0.2108 0.1828 0.1450 0.1050 0.0742 0.0126 0.1556
Bucket 3 0.2860 0.3000 0.2057 0.2034 0.2018 0.1715 0.1424 0.1149 0.1099 0.1878
Bucket 4 0.3000 0.2921 0.3072 0.3101 0.3097 0.3021 0.3239 0.2311 0.2810 0.3016
Bucket 5 0.3709 0.3239 0.3072 0.3000 0.2909 0.2810 0.2311 0.2921 0.3072 0.3016
Title Bucket 1 0.5000 0.4358 0.3438 0.1232 0.1000 0.0778 0.0625 0.0504 0.0493 0.1755
Bucket 2 0.3636 0.2935 0.2826 0.2692 0.1355 0.0997 0.0781 0.0576 0.0560 0.1685
Bucket 3 0.4379 0.4032 0.3704 0.3085 0.2953 0.1464 0.1268 0.0907 0.0897 0.2349
Bucket 4 0.4146 0.3699 0.3201 0.2995 0.1524 0.1506 0.1314 0.1259 0.1184 0.2182
Bucket 5 0.5000 0.4730 0.4545 0.4489 0.4389 0.4231 0.4213 0.3905 0.2822 0.4095
Table 5: DP scores for Slogan and Title Buckets. Values are subtracted by 0.5 to see values under 0.5 easily.
F.2 Discriminative Power of DSAI-Generated Features
To evaluate the reliability of features from DSAI, we organized them into five buckets along the prominence intensity spectrum. From each bucket, we randomly sampled 10 features to annotate the entire dataset with Conform/Not Conform labels when applicable, and marked Not-Applicable when features were inapplicable. For example, we marked "Avoid focusing solely on financial aspects when describing the business model" as Not-Applicable for data points that did not have any business model descriptions.
The analysis revealed that features in the lowest-prominence bucket (Bucket 1) exhibited lower discriminative power across both slogan and title datasets, while features in higher-prominence buckets demonstrated greater reliability (Table 5). This highlights the importance of implementing prominence-based thresholds to filter out less effective features. Notably, none of the DSAI-generated features had a DP score below zero, indicating that no feature mischaracterized negative data traits as positive. This result contrasts sharply with PosData and MixedData) which produced some features with negative DP scores.
These results validate that DSAI systematically filters out less impactful or unreliable features with the use of prominence-based thresholds, enabling more precise and reliable data annotation.
Appendix G Prominence Intensity and Occurrence Analysis
@@
G.1 Influence of Feature Occurrence on Discriminative Power
@@
G.2 Influence of Feature Prominence on Discriminative Power
@@
Appendix H Threshold Analysis on Slogans
@@
H.1 Analysis of Expert-defined Criteria
@@
H.2 Prominence Intensity Threshold Analysis
@@
H.3 Frequency Threshold Analysis
@@
Appendix I Threshold Analysis on Research Titles
Similar to our analysis on slogans (Appendix H), the results suggest that DSAI can effectively capture the nuanced reality of title construction, where practical implementation patterns may differ from theoretical guidelines.
I.1 Analysis of Expert-defined Criteria
@@
I.2 Prominence Intensity Threshold Analysis
@@
I.3 Frequency Threshold Analysis
@@
Appendix J Reliability of Pipeline Operations
In this section, we assess the reliability of our methodology by examining potential errors that could compromise the data-groundedness of the generated features. We specifically focus on cases where incorrect mapping between datapoints and features undermines the accuracy of the feature extraction process.
#1 Perspective Generation
This stage is conceptually error-free by design, as there is no strict requirement for direct correspondence between the generated perspectives and the dataset. While the dataset provides heuristic guidance, the perspectives can be freely generated without compromising the methodology’s integritiy. Therefore, the concept of errors is not applicable to this stage.
#2 Perspective Value Matching Reliability check
Errors can occur when incorrect values are generated, resulting in improper mapping between datapoints and the corresponding perspective-value pairs. To evaluate this, we performed a consistency check on the generated values using LLMs, examining at least three distinct sentences per perspective. The consistency check yielded high accuracy rates exceeding 98% for both research titles and slogans. Specifically, the error rates were 0.042 for slogans (11 out of 258), 0.071 for titles (29 out of 405), and 0.053 for the MIND dataset (15 out of 279). The prompt we used for consistency check can be found in Figure 28.
#3 Value Clustering Reliability check
Errors in this stage occur when cluster labels are incorrectly generated, causing mis-mapping during the conversion from perspective-value to perspective-label forms. This mis-mapping can distort the intended analysis. We conducted the same consistency check as in stage #2, substituting values with labels as input. The accuracy rates exceeded 94% for both titles and slogans.
#4 Verbalization Reliability check
Errors can also arise during the verbalization of perspective-label pairs. To evaluate this, we selected 20 features from each feature bucket generated in Section F.2. We assessed each feature according to two criteria: (1) whether the sentence captures the combined meaning of the perspective and label, and (2) whether the sentence correctly applies directional score by rewriting it in a direct form if $P(\text{positive}|\text{conform})>0.5$, or in an avoidant form if $P(\text{positive}|\text{conform})\leq 0.5$. This evaluation used a detailed prompt requiring True/False responses for each criterion, with returned results in JSON format. The verbalization reliability check demonstrated high accuracy exceeding 98% for both titles and slogans, confirming that the generated sentences faithfully represent the intended meaning. The prompt we used for reliability check can be found in Figure 29.
#5 Prominence-based Selection
This stage is inherently error-free as it operates deterministically through mathematical calculations without involving LLMs.
Appendix K DSAI-Generated Top/Bottom Prominence Features of Expert-Driven Annotation Dataset
@@
Appendix L DSAI-Generated Top/Bottom Prominence Features of Industry Dataset
L.1 Prominence Distribution
@@
L.2 General vs Specific Features
@@
The top and bottom 20 prominence fearures for MIND, SPAM and Reddit dataset are provided in Table 13, 14, 15, respectively.
@@
Figure 12: LLM Annotation prompt
Figure 13: Prompts for direct LLM feature generation.
@@
Figure 23: #1 Perspective Generation Prompt and Outputs
Figure 24: #2 Perspective Value Matching Prompt and Outputs
Figure 25: #3 Clustering Prompt
Figure 26: #3 Clustering Outputs
Figure 27: #4 Verbalization Prompt and Outpus
Figure 28: #2 Perspective Value Matching Reliability Test Prompt
Figure 29: #4 Verbalization Reliability Test Prompt

### v2

@@ 7 Conclusion
8 Limitation
While our results confirm that DSAI can effectively guide LLMs to produce data-grounded features, there are still several limitations to consider. First, the framework’s performance depends heavily on the quality of the underlying LLM. If the model struggles to assign values, cluster them appropriately, or generate coherent perspectives, the outputs can be error-prone or may require significant manual intervention. Second, although GPT-4o was used for annotation and evaluation in some of our experiments, relying on model-based annotations can introduce bias or label noise, potentially affecting overall accuracy. We mitigated this risk by referencing expert-defined criteria and performing manual reviews, but model-based annotations remain a potential source of error.
In addition, DSAI’s current design has thus far been demonstrated primarily on text data, so extending it to other modalities like images, audio, or structured logs may require specialized adaptations or the use of different LLMs. Another limitation involves computational cost and scalability: when evaluating data points across multiple perspectives and then clustering them, running DSAI on very large datasets could become computationally expensive. Employing optimization or sampling strategies might therefore be necessary to maintain efficiency. Finally, despite reducing the reliance on domain experts, some degree of human oversight may still be necessary, particularly in high-stakes environments such as legal or medical settings where interpretability and correctness are paramount.
Appendix A LLM Annotation with Expert-Defined Criteria
@@
A.1 Method
We established a gold standard by manually annotating 10–20 samples per dataset and using expert-defined criteria. After optimizing prompts to maximize alignment with this manual annotation, we used the most aligned prompt to annotate 3,000 data points with GPT-4o, classifying each as "Feature-Present" or "Feature-Absent". This score for each data point was calculated as the total number of conforming criteria. Based on these scores, we selected the top and bottom 600 samples as positive (high-quality text) and negative (low-quality text) samples respectively.
@@ A.1 Method
  Criterion Top Feature-Present (%) Bottom Feature-Present (%)
@@
Table 3: Ratio of Feature-Present in Top and Bottom for Slogan expert-defined criteria.
A.2 Ensuring Distinction Between Positive and Negative Groups
Given that some expert criteria may lack sufficient discriminative power despite their theoretical importance, we conducted an analysis comparing the ratios of "Feature-Present" for each criterion between top- and bottom-ranked samples. Note that a truly discriminative criterion should show high Feature-Present ratio in the top group and low Feature-Present ratio in the bottom group.
  Checklist Item Top Feature-Present (%) Bottom Feature-Present (%)
@@ A.2 Ensuring Distinction Between Positive and Negative Groups
Table 4: Ratio of Feature-Present in Top and Bottom for Title expert-defined criteria.
@@
Table 3 demonstrate significant Feature-Present ratio disparities between the top and bottom groups for most criteria. Several criteria demonstrate maximal distinction (e.g., "Concise but not too simple" (100% vs. 0%)), and some show moderate but still meaningful differentiation (e.g., "Include the brand name" (99% vs. 30%)). "No exaggeration" (99% vs. 91%) and "Direct/Straight-forward" (99% vs. 61%) have smaller gaps.
@@
Table 4 shows even stronger distinction between the top and bottom groups. Multiple criteria (e.g., "Simple format" "Direct", "Concise and precise") achieve 100% Feature-Present ratio in the top group compared to as low as 0–3% Feature-Present ratio in the bottom group, indicating their high predictive value for title quality. The large differences across all criteria confirm their strong discriminative power.
Appendix B Implementation Details and Cost Analysis
@@
In Stage #1 Perspective Generation, we generated 50 perspectives at per forwarding step across three steps, iteratively concatenating each step’s output with the few-shot example in the prompt for the next step. For Stage #2 Perspective-Value Matching, each forwarding step processed one sentence and three perspectives as input. Stage #3 Perspective-Oriented Value Clustering used all generated values per perspective in order to effectively cluster shared characteristics. In Stage #4 Verbalization, we transformed each perspective-label pair into a compact sentence (criterion). The prompts used in this process will be released through github repository.
@@
Appendix C Prominence Intensity and Occurrence Analysis
@@
C.1 Influence of Feature Occurrence on Discriminative Power
@@
C.2 Influence of Feature Prominence on Discriminative Power
@@
Appendix D Threshold Analysis on Slogans
@@
D.1 Analysis of Expert-defined Criteria
@@
D.2 Prominence Intensity Threshold Analysis
@@
D.3 Frequency Threshold Analysis
@@
Appendix E Threshold Analysis on Research Titles
Similar to our analysis on slogans (Appendix D), the results suggest that DSAI can effectively capture the nuanced reality of title construction, where practical implementation patterns may differ from theoretical guidelines.
E.1 Analysis of Expert-defined Criteria
@@
E.2 Prominence Intensity Threshold Analysis
@@
E.3 Frequency Threshold Analysis
@@
Appendix F DSAI-Generated Top/Bottom Prominence Features of Expert-Driven Annotation Dataset
@@
Appendix G DSAI-Generated Top/Bottom Prominence Features of Industry Dataset
G.1 Prominence Distribution
@@
G.2 General vs Specific Features
@@
The top and bottom 20 prominence fearures for MIND, SPAM and Reddit dataset are provided in Table 12, 13, 14, respectively.